To make a 10 pound mix of M&M's and peanuts worth $3.50 a pound, how many pounds of peanuts worth $3 per pound should be mixed with M&M's worth $4 per pound?
I have tried;
x=$3x+$4x
x=7
7/10=3 1/2
10(7/10)=10*3 1/2
7=30*2
= 60/7
not even sure now how I got my answer, but I know it is wrong.

Found 2 solutions by scott8148, Alan3354:
Answer by scott8148(6628)

x=peanuts, so 10-x=M&M

3(x) + 4(10-x) = 3.5(10)

3x + 40 - 4x = 35

-x = -5

x = 5

Answer by Alan3354(69443)

$3.50 is the center between $3 and $4, so it's 50% of each.